# Quantum variational calculus on a lattice

### v1

@@ 4.1. EL equations and translation symmetries for $X=\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$
Comparing the expressions to what was found in the continuum case in Example 2.6, we see that these are similar. In the classical limit, where we expect $\partial_{t}\phi$, $\partial_{t^{-1}}\phi$ to correspond to ±$\pm\partial_{0}\phi$, and the spatial derivatives $\partial_{b}\phi$, $\partial_{b^{-1}}\phi$ to ±$\pm\partial_{i}\phi$, we see that these expressions recover exactly the classical forms for $E$ and $P_{i}$ for the Euclidean metric $g_{00}=g_{ii}=1$. The extra 1/21/21 / 2 is there in order to account for the doubling of the derivatives.
@@
$T_{ab}=-\left(\frac{1}{2}u_{a}u_{b}+\frac{1}{4}u_{ab}u_{b}+{\delta^{a^{-1}}_{b}\over\eta^{a}}L\right)$
@@
$E$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}t}[\phi]-R_{t^{-1}}T_{tt}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=-{1\over 2}\int_{\mathbb{Z}^{m-1}}\left((\partial_{t}\phi)(\partial_{t^{-1}}\phi)-{1\over 2}\sum_{b\in\mathcal{C}_{m-1}}(\partial_{b}\phi)^{2}-{m^{2}}\phi^{2}\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}},$
$P_{b}$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}b}[\phi]-R_{t^{-1}}T_{tb}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=-{1\over 2}\int_{\mathbb{Z}^{m-1}}\partial_{t^{-1}}\phi\left(\partial_{b}\phi-\partial_{b^{-1}}\phi\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}}.$
@@ 5. Lagrangians on lattices with background metrics and gauge fields
This section is dedicated to the computation of the EL equations and conserved charges for complex fields with a background non-Euclidean metric and/or a $U(1)$-gauge field to which the field couples. We still work on the base being an Abelian group $X$, but some results can be applied to any digraph using the quantum Riemannian geometry (QRG) formalism [2] and gauge theory as in [21]. We already explained at the start of Section 4 how we work with the differential exterior algebra $(\Omega,\mathrm{d})$ on the algebra $A=C(X)$ corresponding to the Cayley graph induced by a set $\mathcal{C}$ of generators. We only use the group structure to provide convenient coordinates and the left invariant basis 1-forms $\{e^{a}\}$.
@@
$\nabla\omega=\theta\otimes_{A}\omega-\sigma(\omega\otimes_{A}\theta)+\beta(\omega)$
@@
for $\phi,\psi\in A$. Here $\Delta=\delta\mathrm{d}$ where the codifferential or divergence of a 1-form is $\delta=-{1\over 2}(\,,\,)\nabla$. In both formulae, we have introduced a factor $-{1\over 2}$ compared to the canonical QRG normalisations in order to match physics conventions. An integral $\int\colon A\to\mathbb{C}$ (in our case given by sum over $X$ with a measure $\mu$) is said to be divergence compatible if $\int\circ\delta=0$.
@@ 5.1. Scalar field theory on $X$ with a generic metric
(2) In this case integration with measure $\mu$ is divergence-compatible w.r.t. $\delta$ iff
@@
(1) Explicitly, from the form of $\sigma$, $\beta$ and $\theta$
@@
$\delta e^{a}=-{1\over 2}\left(\lambda^{a^{-1}}-\sum_{b\in\mathcal{C}}\lambda^{b}\sigma^{aa^{-1}}{}_{b}\right).$
@@
$-2\int\delta(e^{a})$ $=\int_{X}\left(\lambda^{a^{-1}}-\sum_{b\in\mathcal{C}}\lambda^{b}\sigma^{aa^{-1}}{}_{b}\right)\mu=\sum_{g\in X}\left(\mu(g)\lambda_{g\to ga^{-1}}-\sum_{b\in\mathcal{C}}\mu(g)\lambda_{g\to gb}\sigma^{ab}{}_{c}(g)\right)$
$=\sum_{g\in X}\left(\mu(ga)\lambda_{ga\to g}-\sum_{b\in\mathcal{C}}\mu(g)\lambda_{g\to gb}\sigma^{aa^{-1}}{}_{b}(g)\right)$
for all $g$. Vanishing of this gives is the condition stated.
@@
For a connected calculus the condition in part (2) uniquely determines $\mu$, if it exists, up to an overall constant from the metric coefficients $\lambda_{a}$ and the braiding $\sigma$. This is a graph version of the way the Riemannian measure of integration is determined by the metric.
@@
Equation (20) extremizes the action
@@
(2) In this case integration with measure $\mu$ is divergence-compatible wrt $\delta=(\ ,\ )\nabla$, i.e. $\int\circ\delta=0$, iff

### v2

@@ 4.1. EL equations and translation symmetries for $X=\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$
Comparing the expressions to what was found in the continuum case in Example 2.6, we see that these are similar. In the classical limit, where we expect $\partial_{t}\phi$, $\partial_{t^{-1}}\phi$ to correspond to $\pm\partial_{0}\phi$, and the spatial derivatives $\partial_{b}\phi$, $\partial_{b^{-1}}\phi$ to $\pm\partial_{i}\phi$, we see that these expressions recover exactly the classical forms for $E$ and $P_{i}$ for the Euclidean metric $g_{00}=g_{ii}=1$. The extra $1/2$ is there in order to account for the doubling of the derivatives.
4.2. Scalar field theory in the (1+1)-dimensional case
We focus now on (1+1)-dimensional case of Euclidean field theory, where the Klein-Gordon equation $(\Delta+m^{2})\phi(t,x)=0$ is solved by the plane wave solutions
$\phi(t,x)=A\cos(\omega t+\kappa x)$
satisfying as well known dispersion relation for a lattice
(20) $m^{2}=2(2-\cos(\omega)-\cos(\kappa)).$
as in [23, 25, 26]. This is normally justified from its appearance in poles in the propagator after Fourier transform to $\omega,\kappa$ as the Fourier conjugate variables, whereas our considerations here will be in line with the continuum version from classical variational calculus in Example 2.6, for which we have the required limit $m^{2}\simeq\omega^{2}+\kappa^{2}$ for $\omega,\kappa\ll 1$. Note that in Section 3.3, we saw that a classical particle $q$ in a quadratic potential $V(q)=\frac{1}{2}m\omega^{2}q^{2}$ only allowed for plane-wave solutions for $\omega<2$, with a different type of numerical solutions appearing for $\omega>2$. Here, we similarly find that the mass $m$ of the field for a plane wave (now playing the role of the frequency $\omega$ in the previous example) is also bounded as $m\leq 2\sqrt{2}$ as $\cos(\omega),\cos(\kappa)\in[-1,1]$.
The relation (20) can be inverted to give
(21) $\omega(\kappa)=\arccos\left(2-\cos(\kappa)-{m^{2}\over 2}\right)$
which is defined for all $\kappa\in[-\pi,\pi]$ that satisfy $\cos(\kappa)\in\left[1-{m^{2}\over 2},3-{m^{2}\over 2}\right]\cap\left[-1,1\right]$. For example for the values of $m=0,2,2\sqrt{2}$ we find $\cos(\kappa)=1$, $\cos(\kappa)\in[-1,1]$ and $\cos(\kappa)=-1$ respectively. The dispersion relation is illustrated in Figure 2 at essentially these values of $m$, where it is also compared with the continuum counter part shown dashed. We see that the results match for small masses, but are very different as the fields become more massive. In particular, for massless modes we have $\kappa=0$, for modes with mass $m=2$ we see that all $\kappa$ are allowed, and as we approach $m=2\sqrt{2}$ we have that only the modes close to $\kappa=-\pi,\pi$ propagate. We also observe that the dispersion relation (20) has an interesting symmetry
$m\mapsto m^{\prime}=\sqrt{8-m^{2}},\quad\omega\mapsto\omega^{\prime}=\pi-\omega,\quad\kappa\mapsto\kappa^{\prime}=\pi-\kappa$
with the fixed point at $m^{\prime}=m=2$. For example, $m\to 0$, $\omega(\kappa)$ ‘collapses’ to the point $\omega=\kappa=0$ in Figure 2, while in the limit $m\to 2\sqrt{2}$, the dispersion relation mirrors this behaviour but towards the point $\omega=\pi$, $\kappa=\pi$. From $\omega(\kappa)$, we find that the phase and group velocities are
$v_{ph}=\frac{\omega(\kappa)}{\kappa},\quad\quad v_{gr}=\frac{\partial\omega(\kappa)}{\partial\kappa}=-\frac{\sin(\kappa)}{\sqrt{1-\left(2-\cos(\kappa)-\frac{m^{2}}{2}\right)^{2}}},$
as illustrated in Figure 3.
Figure 2. Plots of the dispersion relation, energy and momentum densities at masses $m=0.9,1.7,2,2.25,2.67$ and comparison with the continuum counterparts in the Euclidean case dashed. The $m>2$ values are the $m^{\prime}=\sqrt{8-m^{2}}$ counterparts of the $m<2$ values but with 0.01-0.01 offset to separate curves in the last plot. For small masses, the lattice and continuum plots almost match. At $m=2$, waves with all $\kappa\in[-\pi,\pi]$ propagate in the lattice. As we approach $m=2\sqrt{2}$, waves closer to $\kappa=0$ stop propagating first, until eventually at $m=2\sqrt{2}$, only waves with $\kappa=-\pi,\pi$ propagate. The lower right plot shows the energy density against the momentum density for $A=\sqrt{2/\langle\mathcal{E}_{dens}[\phi]\rangle}$ as in Equation (24).
Regarding the energy and momentum densities, we find that they are given by
$\mathcal{E}_{dens}[\phi]$ $=-{1\over 2}(\partial_{t}\phi)(\partial_{t^{-1}}\phi)-{1\over 4}(\partial_{x}\phi)^{2}+{1\over 2}m^{2}\phi^{2}$
$=\frac{A^{2}}{8}\left(4\cos^{2}(\kappa x+t\omega)-\cos(2\kappa(x-1)+2t\omega)-\cos(2(\kappa x+\kappa+t\omega))-2\cos(2\omega)\right)$
$\mathcal{P}_{x,dens}[\phi]$ $={1\over 2}(\partial_{t^{-1}}\phi)\left(\partial_{x^{-1}}\phi-\partial_{x}\phi\right)=A^{2}\sin(\kappa)\sin(\omega)\sin^{2}\left(\kappa x+\frac{\kappa}{2}+t\omega-\frac{\omega}{2}\right)$
As we saw in Example 2.6, these densities do not make it directly clear that the energy $E[\phi]$ and momenta $P_{x}[\phi]$ are time independent. In the continuum, we took the spatial average over a period $2\pi/\kappa$ and found that it is time independent for both quantities, signifying that the total energy and momentum will also be. In the discrete case, taking the average over a period $2\pi/\kappa$ is not natural as the period is not an integer in general. Instead, we average over the whole $\mathbb{Z}$ line to find
$\langle\mathcal{E}_{dens}[\phi]\rangle=\lim_{N\to\infty}\frac{1}{2N}\sum^{x=N}_{x=-N}\mathcal{E}_{dens}[\phi]=\frac{A^{2}}{2}\sin^{2}(\omega),$
$\langle\mathcal{P}_{x,dens}[\phi]\rangle=\frac{A^{2}}{2}\sin(\kappa)\sin(\omega).$
In both cases, the classical limit where $\omega$ and $\kappa$ are small recovers the expressions from Example 2.6 in the Euclidean case as a useful check on our reasoning. As expected, we found that both the energy and momentum densities are bounded as
(22) $\langle\mathcal{E}_{dens}[\phi]\rangle\in\left[0,\frac{A^{2}}{2}\right]\quad\quad\langle\mathcal{P}_{x,dens}[\phi]\rangle\in\left[-\frac{A^{2}}{2},\frac{A^{2}}{2}\right]$
due to lattice effects. These densities and their comparison with the classical expressions are also shown in Figure 2. The symmetry between $m$ and $m^{\prime}=\sqrt{8-m^{2}}$ is also present in the energy and momentum density, where the plots for $m=0.9,1.7$ centred around $\kappa$ are mirrored by plots for $m^{\prime}=2.25,2.67$ now centered around $\kappa=\pi$. We did not take exactly $\sqrt{8-m^{2}}$ so that the curves would not be exactly on top of each other in the energy-momentum plot.
Figure 3. Plots for the phase and group velocity for different masses as in Figure 2.
As in Example 2.6, we find a relation between the energy and momentum densities through the dispersion relation (20). We start by writing $\cos(\omega)=s_{\omega}\sqrt{1-\sin^{2}(\omega)}$ and by $\cos(\kappa)=s_{\kappa}\sqrt{1-\sin^{2}(\kappa)}$ for $s_{\omega},\,s_{\kappa}\in\{\pm 1\}$. Using the formulas for the energy and momentum density above to write $\sin^{2}(\omega)$ and $\sin^{2}(\kappa)$ in terms of $\langle\mathcal{E}_{dens}[\phi]\rangle$, $\langle\mathcal{P}_{x,dens}[\phi]\rangle$, and reordering the terms results in the energy and momentum densities relation
(23) $\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}$ $=\frac{A^{2}}{2}\langle\mathcal{E}_{dens}[\phi]\rangle\left(1-\left(2-\frac{m^{2}}{2}-s_{\omega}\sqrt{1-\frac{2\langle\mathcal{E}_{dens}[\phi]\rangle}{A^{2}}}\right)^{2}\right)$
$\approx\left(-1+\frac{m^{2}}{2}\right)\langle\mathcal{E}_{dens}[\phi]\rangle^{2}+\frac{A^{2}}{2}\langle\mathcal{E}_{dens}[\phi]\rangle m^{2}\left(1-\frac{m^{2}}{4}\right),$
where the approximation was made for $\omega\ll 1$ (i.e. $\langle\mathcal{E}_{dens}[\phi]\rangle\ll 1$) much smaller than the lattice spacing, and thus $s_{\omega}=1$ since $\cos(\omega)>0$ in this case. This clearly deforms the classical relation $\langle\mathcal{E}_{dens}[\phi]\rangle^{2}+\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}=(A^{2}/2)\langle\mathcal{E}_{dens}[\phi]\rangle^{2}\,m^{2}$ from Example 2.6 in the Euclidean case.
Following the strategy in Example 2.6, we set $A=\sqrt{2/\langle\mathcal{E}_{dens}[\phi]\rangle}$ to have a field $\phi$ which has a ‘density of one particle per unit volume’. In this case, the dispersion relation (20) reads
$\frac{m^{2}}{2}=2-s_{\omega}\sqrt{1-\langle\mathcal{E}_{dens}[\phi]\rangle^{2}}-s_{\kappa}\sqrt{1-\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}},$
with energy density solutions
(24) $\langle\mathcal{E}_{dens}[\phi]\rangle_{s_{\kappa}=\pm 1}=\sqrt{1-\left(2-\frac{m^{2}}{2}-s_{\kappa}\sqrt{1-\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}}\right)^{2}}.$
This relation results in the energy against momentum plot in Figure 2, where the $s_{\kappa}=+1$ (small $\kappa\ll 1$) solution is comparable to the classical case $\langle\mathcal{E}_{x,dens}[\phi]\rangle=\sqrt{m^{2}-\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}}$ shown dashed. The solutions (24) again display the $m,\,m^{\prime}=\sqrt{8-m^{2}}$ symmetry in that $\langle\mathcal{E}_{dens}[\phi]\rangle_{\pm}$ for $m$ correspond to $\langle\mathcal{E}_{dens}[\phi]\rangle_{\mp}$ for $m^{\prime}$.
@@ 4.2. Scalar field theory in the (1+1)-dimensional case
$T_{ab}=-\left(\frac{1}{2}u_{a}u_{b}+\frac{1}{4}u_{ab}u_{b}+\frac{\delta^{a^{-1}}_{b}}{\eta^{a}}L\right)$
@@
$E[\phi]$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}t}[\phi]-R_{t^{-1}}T_{tt}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=\int_{\mathbb{Z}^{m-1}}{1\over 2}\left(-(\partial_{t}\phi)(\partial_{t^{-1}}\phi)+{1\over 2}\sum_{b\in\mathcal{C}_{m-1}}(\partial_{b}\phi)^{2}+{m^{2}}\phi^{2}\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}},$
$P_{b}[\phi]$ $=\int_{\mathbb{Z}^{m-1}}(T_{t^{-1}b}[\phi]-R_{t^{-1}}T_{tb}[\phi])\mathrm{Vol}_{\mathbb{Z}^{m-1}}$
$=\int_{\mathbb{Z}^{m-1}}{1\over 2}\partial_{t^{-1}}\phi\left(\partial_{b^{-1}}\phi-\partial_{b}\phi\right)\mathrm{Vol}_{\mathbb{Z}^{m-1}}.$
@@
Figure 4. Dipersion relation, energy density and momentum densities in the Minkowski case for the masses $m=0.5,\sqrt{2},1.9$. At lower masses, $\kappa=0$ is a local minimum of the energy density, while it is a local maximum for waves with higher masses. In the lower right, the energy density is plotted against the momentum density following Equation (28).
For the (1+1)-dimensional Minkowski case we now consider the plane wave solutions $\phi(t,x)=A\cos(\omega t-\kappa x)$ of the Klein-Gordon equation. In this case, the dispersion relation reads instead
(26) $m^{2}=2(\cos(\kappa)-\cos(\omega))$
as also found in [14]. This recovers the classical limit $m^{2}=\omega^{2}-\kappa^{2}$ when $\omega,\kappa\ll 1$, i.e. much smaller than the lattice spacing. We find that in this case the mass is bounded as $m<2$ for physical states, and find that the frequency behaves as
$\omega(\kappa)=\arccos\left(\cos(\kappa)-{m^{2}\over 2}\right)$
with now the wavelength having to satisfy $\cos(\kappa)\in\left[-1+{m^{2}\over 2},1\right]$. Contrary to the Euclidean case, we find that massless modes can propagate with any wavelength $\kappa$, while the only allowed mode at $m=2$ is the constant one with $\kappa=0$. The dispersion relation is illustrated in Figure 4 for $m=0.5,\sqrt{2},1.9$, where it is also compared with the classical counter part. The phase velocity is again given by $v_{ph}=\omega(\kappa)/\kappa$ whereas the group velocity now reads
$v_{gr}=\frac{\partial\omega(\kappa)}{\partial\kappa}=\frac{\sin(\kappa)}{\sqrt{1-\left(\cos(\kappa)-\frac{m^{2}}{2}\right)^{2}}},$
as illustrated in Figure 5 for the same values of $m^{2}$ as before, where they are also compared with the classical counter part.
The analysis of the energy and momentum densities goes as in the Euclidean case, where we now find
$\mathcal{E}_{dens}[\phi]$ $=\frac{A^{2}}{2}\left(\sin^{2}(\omega)-\sin^{2}(\kappa)\cos(2t\omega-2\kappa x)\right),$
$\mathcal{P}_{x,dens}[\phi]$ $=-A^{2}\sin(\kappa)\sin(\omega)\sin^{2}\left(\left(t-\frac{1}{2}\right)\omega-\frac{1}{2}\kappa(2x+1)\right)$
and the averages
$\langle\mathcal{E}_{dens}[\phi]\rangle=\frac{A^{2}}{2}\sin^{2}(\omega),\quad\quad\langle\mathcal{P}_{x,dens}[\phi]\rangle=-\frac{A^{2}}{2}\sin(\kappa)\sin(\omega),$
as before, here with a minus in the momentum density due to the metric signature. As expected, these are time independent and tend to the continuum values for small $\omega,\kappa$. They are also plotted in Figure 4. We see that the profile of the energy density changes depending on whether $m<\sqrt{2}$, $m=\sqrt{2}$ or $m>\sqrt{2}$. In the first case, there are two local maxima and 3 local minima, one at $\kappa=0$ and two at the edges of the domain. For $m>\sqrt{2}$, the point at $\kappa=0$ is a local maximum. Comparing these with the plot of the momentum density, we find that these energy minima, either at $\kappa=0$ or at the edges of the $\kappa$-domain, always correspond to plane waves with zero momentum density.
Figure 5. Phase and group velocities in the Minkwoski case for $m=0.5,\sqrt{2},1.9$ and comparison with their classical counterparts shown dashed.
As in the Euclidean case, we can relate the energy and momentum density using the dispersion relation (26). In this case we find
(27) $\langle\mathcal{P}_{x,dens}\rangle^{2}$ $=\frac{A^{2}}{2}\langle\mathcal{E}_{dens}\rangle\left(1-\left(\frac{m^{2}}{2}+s_{\omega}\sqrt{1-\frac{2\langle\mathcal{E}_{dens}\rangle}{A^{2}}}\right)^{2}\right)$
$\approx\left(1+\frac{m^{2}}{2}\right)\langle\mathcal{E}_{dens}\rangle^{2}-\left(1+{m^{2}\over 4}\right)\frac{A^{2}}{2}\langle\mathcal{E}_{dens}\rangle\,m^{2}$
where the approximation is made for $\omega\ll 1$, which results in a deformation of the classical relation $\langle\mathcal{E}_{dens}\rangle^{2}-\langle\mathcal{P}_{x,dens}\rangle^{2}=(A^{2}/2)\langle\mathcal{E}_{dens}\rangle\,m^{2}$ from Example 2.6 for the Minkowski case. We can again set $A=\sqrt{2/\langle\mathcal{E}_{dens}\rangle}$, and following similar steps to the Euclidean case, we find the energy density solutions
(28) $\langle\mathcal{E}_{dens}[\phi]\rangle_{s_{\kappa}=\pm 1}=\sqrt{1-\left(\frac{m^{2}}{2}-s_{\kappa}\sqrt{1-\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}}\right)^{2}}$
as illustrated in Figure 4 against their classical counterparts. For small $\kappa$, the relevant solution that deforms the classical case is with $s_{\kappa}=+1$ as in the introduction.
@@ 5. Lagrangians on lattices with background metrics and gauge fields
This section is dedicated to the computation of the EL equations and conserved charges for complex fields with a background non-Euclidean metric and/or a $U(1)$-gauge field to which the field couples. We still work on the base being an Abelian group $X$, but some results can be applied to any digraph using the quantum Riemannian geometry (QRG) formalism [2] and gauge theory as in [22]. We already explained at the start of Section 4 how we work with the differential exterior algebra $(\Omega,\mathrm{d})$ on $A=C(X)$ corresponding to the Cayley graph induced by a set $\mathcal{C}$ of generators. We only use the group structure to provide convenient coordinates and the left invariant basis 1-forms $\{e^{a}\}$.
@@
$\nabla\omega=\theta\otimes_{A}\omega-\sigma(\omega\otimes_{A}\theta)+\beta(\omega),$
@@
[…]
@@ 5.1. Scalar field theory on $X$ with a generic metric
(2) In this case, integration with measure $\mu$ is divergence-compatible w.r.t. $\delta$ iff
@@
(1) Explicitly, from the form of $\sigma$, $\beta$ and $\theta$,
@@
[…]
@@
[…]
for all $g$. Vanishing of this gives the condition stated.
@@
For a connected calculus, the condition in part (2) uniquely determines $\mu$, if it exists, up to an overall constant from the metric coefficients $\lambda_{a}$ and the braiding $\sigma$. This is a graph version of the way the Riemannian measure of integration is determined by the metric.
@@
Equation (33) extremises the action
@@
(2) In this case, integration with measure $\mu$ is divergence-compatible wrt $\delta=(\ ,\ )\nabla$, i.e. $\int\circ\delta=0$, iff